# Video-Language Critic: Transferable Reward Functions for Language-Conditioned Robotics

### v2

@@ 4 Experiments
We evaluate the accuracy and effectiveness of the learned video-language rewards on simulated robotic manipulation tasks from the Meta-World benchmark [11]. We evaluate VLC’s ability to inform successful policy training in three settings of increasing difficulty. First, we assess the ability of our model to jointly represent several robotic tasks with a single language-conditioned prediction network in Section 4.1. Second, we test our models’ ability to generalize to unseen Meta-World tasks with the help of vision-language pretraining as well as extrapolation from training tasks in Section 4.2. In Section 4.3, we demonstrate our method’s out-of-domain transfer ability: VLC is used to learn an embodiment-agnostic reward function for any language-conditioned manipulation task by observing a variety of robot actors from Open X-Embodiment [10], a large dataset collected from a variety of real-world robots in different environments. We further report comparisons to prior work, both quantitatively and qualitatively, in Section 4.4, and demonstrate VLC’s effectiveness in planning with a known dynamical model in Section 4.5. Finally, in Appendix E, we investigate the model’s robustness to visual variations (brightness shift and image noise).
@@
We perform hyperparameter selection, ablation studies, and finalize all training details on data from VLMBench [40], a second robotic benchmark. This avoids Meta-World specific tuning and allows for a fairer comparison with prior work. Details on this dataset and our model ablations are included in Appendix A. We observe the same hyperparameters perform well both on VLMBench videos and in interactive RL policy training on Meta-World, which highlights the method’s applicability across domains. Further VLC training details are reported in Appendix B. To train control policies, we use Soft Actor-Critic (SAC) [41], an off-policy RL algorithm. As the reward function, we use either a sparse task completion reward, equal to $1$ if the task has been solved and 00 otherwise, or a weighted sum of the learned VLC reward and the sparse reward. As the similarity $S_{\theta}(v_{1:t},c)$ predicted by VLC can take arbitrary values, we apply a few normalization operations to aid in the stability of RL optimization [42] (details in Appendix B.2).
@@
We evaluate VLC in RL policy training on robotic manipulation tasks from the Meta-World benchmark. As the focus of this work is the problem of learning a foundation reward model, we keep to a standard single-task policy training setup and condition on full state information as defined by Meta-World (see Appendix C for more details on the environment definitions). Compared to using image-only observations, this reduces training time and computational cost while allowing us to demonstrate the effectiveness of our reward function. We consider the subset of tasks that can be reliably solved using the dense Meta-World rewards, which are manually specified for each environment based on the full state of the environment. Specifically, we include tasks that can be solved with $\geq$98success within a maximum training length of 800,000 steps. We further split these in half into 12 easy (learned in $<$240,000 steps) and 13 hard tasks (240k – 800k steps).
@@
{tblr}
l c c c & sparse only VLC MW50 VLC Open X
easy (12) 53 $\pm$ 6 65 $\pm$ 5 54 $\pm$ 6
hard (13) 62 $\pm$ 5 91 $\pm$ 3 72 $\pm$ 5
mean 58 $\pm$ 4 78 $\pm$ 3 63 $\pm$ 4
@@ 4.1 Multi-task reward function
The policy training results are shown in Table 4, with learning curves for the hard task set in Fig. 2. To summarize learning speed with a single number, we report success rate of the policy evaluated at the training length at which the manually specified Meta-World reward solves the task to $\geq$98success. VLC trained on MW50 enables improved sample efficiency relative to the sparse reward only, which demonstrates that VLC can sufficiently generalize to trajectories not seen in demonstration data, and can effectively represent task progress for multiple tasks at once. However, a few tasks, such as Handle Press, are learned in so few trials even with sparse reward alone that there is little room for improvement in reward design, and learning is instead bottlenecked by the policy training’s sample efficiency. This is why the biggest gains are obtained for the harder tasks.
@@ 4.2 Task generalization to unseen environments
Success rates of RL policies on these 6 held-out tasks are shown in Table 4.2 (calculated similarly to Table 4), with learning curves in Fig. 3. In most tasks, the addition of VLC rewards improves sample efficiency of RL training despite the tasks having never been seen in reward function training, with average success rates 34 percentage points over the sparse baseline and an average sample efficiency improvement of over 5x.
In addition to task generalization, we further evaluate VLC in out-of-distribution visual conditions in Appendix E.
@@
{tblr}
l c c c c c c LOReL RoboCLIP LIV Voltron R3M VLC (ours) sparse only
easy (2) 57 $\pm$ 21 46 $\pm$ 26 30 $\pm$ 10 18 $\pm$ 10 22 $\pm$ 11 58 $\pm$ 22 39 $\pm$ 12
hard (4) 60 $\pm$ 16 51 $\pm$ 11 63 $\pm$ 15 35 $\pm$ 11 24 $\pm$ 14 79 $\pm$ 18 42 $\pm$ 10
mean 59 $\pm$ 19 50 $\pm$ 12 39 $\pm$ 11 29 $\pm$ 19 16 $\pm$ 10 72 $\pm$ 11 41 $\pm$ 18
@@ 4.3 Embodiment generalization to unseen domains
The advantage of our method, and pretraining a reward function in general, is that no data collection on the target robot and in the target environment is required. To demonstrate this, we train VLC on cross-embodiment data from Open X-Embodiment [10]. We use the language-annotated subset, with a total of 698,000 episodes of diverse tasks filmed in various real-world robotic labs. Although some of this data does feature the Sawyer robot used in Meta-World simulations, this is only a marginal subset of 0.33% of the language-annotated videos. Moreover, the domain gap remains significant due to real-world variations in objects, backgrounds, lighting conditions, task instances and instruction formats, as well as the embodiment gap between the simulated and the real robots.
We successfully train policies (see Table 4 and Fig. 2) using Open X trained models despite a significant domain gap, highlighting the generalizability of large-scale vision-language training. We obtain an average 2.1x sample efficiency gain over the sparse reward in tasks solved by both rewards (sample efficiency is ill-defined if either does not solve the task), with particularly large improvements in Handle Pull Side (7x), Reach Wall (7x) and Slide Plate (6x) and an average 5 percentage point success rate increase across all 25 tasks despite misrepresenting a few tasks. Note that unlike RT-X [10], our method does not use action labels, and remains equally applicable on observation-only data.
@@ 4.4 Comparison to prior work
To validate the benefits of VLC, we compare its performance to prior language-conditioned reward models LOReL [15], RoboCLIP [16], LIV [17], Voltron [19] and R3M [18], each fine-tuned on MW40. Training and implementation details are included in Appendix D.
We find VLC’s combination of cross-entropy and the sequential ranking objective as well as full video conditioning to produce more informative reward predictions than existing methods, as shown by faster policy training on average in Table 4.2 and Fig. 3. For increased statistical significance, we use 5 additional random seeds (10 total) for VLC and the two strongest baselines: RoboCLIP and LOReL. Moreover, on qualitative inspection of the shape of the predicted rewards (Appendix F), we find VLC’s outputs to better distinguish successes from failures compared to either RoboCLIP or LOReL. Thanks to its broad coverage of execution history and the sequential loss term, VLC also produces rewards that more smoothly increase over successful episodes than either prior method.
@@ 4.5 Model-based evaluation
We find VLC to generalize well to these tasks zero-shot, with 80% mean accuracy on held-out tasks, outperforming RoboCLIP (46%) and LOReL (36%). A breakdown per task is shown in Appendix G.
@@ 5 Conclusion
We proposed Video-Language Critic (VLC), a method for training a foundation reward model for vision-language manipulation. In particular, we train our critic using contrastive video-language alignment and a ranking loss encouraging monotonic increases for successful trajectories. Our model predicts a language-conditioned state-value function conditioned on only a history of image observations, and can therefore readily be scaled to leverage external observation data from other actors. VLC can be used for various downstream tasks, such as model-free and model-based reinforcement learning. Further, we experimentally validated its usefulness as a reward function in out-of-distribution tasks (unseen, held-out environments in the same domain) and out-of-domain tasks (an unseen environment and embodiment in a different domain). Our experiments on Meta-World demonstrated improved results compared to 5 prior methods and a sparse reward baseline, with success rate increases of 12 and 34 percentage units, respectively, and sample efficiency gains of 65% and 300%, respectively. Unlike methods based on predicting expert actions in a given embodiment, VLC remains applicable in the absence of action labels. Therefore, it can be extended to equally learn from videos of humans in future work.
Acknowledgments
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI, and grants 345604, 352986, 358246, 359207), EU Horizon 2020 (European Network of AI Excellence Centres ELISE, grants 951847, 101016775 and NextGenerationEU) and UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1). We acknowledge the computational resources provided by the Aalto Science-IT Project from Computer Science IT, and CSC for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through Finland.
@@ Acknowledgments
Appendix A Model ablations
@@ A.1 VLMBench dataset
We use the VLMBench manipulation task suite to develop and validate our method without any Meta-World specific tuning. For this purpose, we collect 2700 video demonstrations and 1600 failure cases from variations of the picking task – covering different object shapes, sizes, colors and relative positions, as well as distractor objects. The natural language instructions match this diversity in task variants, such as Grasp the cylinder or Grasp the cyan object, and require distinguishing relevant objects from distractors with either absolute (color, shape) or relative (size, position) properties, such as the the larger or the front object. For more details on the benchmark, see Zheng et al. [40].
We use these VLMBench videos to validate VLC design decisions, but defining a single informative metric on the dataset of video-caption pairs $(v^{i},c^{i})$, $i=1,...,N$, is difficult. Test loss, video retrieval metrics such as mean recall, or classification metrics such as area under the ROC curve do not correspond well to the models’ ability to model task progress. The main difficulty is that part of the caption-to-video matching task can be solved by simply connecting objects referred to in the caption to objects present in the scene, without considering temporal information or actual task success.
@@ A.2 Ablation results
We compare two temporal aggregation methods as proposed by Luo et al. [26]: the sequence transformer and the tight-type transformer. The sequence transformer aggregates the sequence of image features $[\text{ViT}(v_{1}),\text{ViT}(v_{2}),...,\text{ViT}(v_{T})]$ into a single embedding vector $S_{\theta}(v_{1:T})$, which it then compares to the caption embedding $\text{TextEnc}(c)$ with cosine similarity. The tight-type transformer, on the other hand, includes the caption embedding as an additional input to the temporal aggregator $S_{\theta}(v_{1:T},c)$, as shown in Fig. 1.
We report the results of our ablation study in Table A.2. In addition to the choice of architecture, we observe performance gains from adding image augmentations from the Albumentations library [43], by sampling frames randomly from uniform intervals instead of deterministic uniform sampling, the addition of the sequence ranking term, as well as considering failure examples only as negatives in the contrastive objective, and report results using these settings in Section 4.
@@
{tblr}
c | l | c | c| c Architecture $\alpha$ Data augmentations Failures Accuracy (%)
@@
Appendix B Training details
B.1 Reward training

### v3

@@ 4 Experiments
We evaluate the accuracy and effectiveness of the learned video-language rewards on simulated robotic manipulation tasks from the Meta-World benchmark (Yu et al., 2019). We evaluate VLC’s ability to inform successful policy training in three settings of increasing difficulty. First, we assess the ability of our model to jointly represent several robotic tasks with a single language-conditioned prediction network in Section 4.1. Second, we test our models’ ability to generalize to unseen Meta-World tasks with the help of vision-language pretraining as well as extrapolation from training tasks in Section 4.2. In Section 4.3, we demonstrate our method’s out-of-domain transfer ability: VLC is used to learn an embodiment-agnostic reward function for any language-conditioned manipulation task by observing a variety of robot actors from Open X-Embodiment (Open X-Embodiment Collaboration et al., 2023), a large dataset collected from a variety of real-world robots in different environments. We further report comparisons to prior work, both quantitatively and qualitatively, in Section 4.4, and demonstrate VLC’s effectiveness in planning with a known dynamical model in Section 4.5. Finally, in Appendix F, we investigate the model’s robustness to visual variations (brightness shift and image noise).
@@
We perform hyperparameter selection, ablation studies, and finalize all training details on data from VLMbench (Zheng et al., 2022), a second robotic benchmark. This avoids Meta-World specific tuning and allows for a fairer comparison with prior work. Details on this dataset and our model ablations are included in Appendix A. We observe the same hyperparameters perform well both on VLMbench videos and in interactive RL policy training on Meta-World, which highlights the method’s applicability across domains. Further VLC training details are reported in Appendix C. To train control policies, we use Soft Actor-Critic (SAC) (Haarnoja et al., 2018), an off-policy RL algorithm. As the reward function, we use either a sparse task completion reward, equal to $1$ if the task has been solved and 0 otherwise, or a weighted sum of the learned VLC reward and the sparse reward. As the similarity $S_{\theta}(v_{1:t},c)$ predicted by VLC can take arbitrary values, we apply a few normalization operations to aid in the stability of RL optimization (van Hasselt et al., 2016) (details in Appendix C.2).
@@
We evaluate VLC in RL policy training on robotic manipulation tasks from the Meta-World benchmark. As the focus of this work is the problem of learning a foundation reward model, we keep to a standard single-task policy training setup and condition on full state information as defined by Meta-World (see Appendix D for more details on the environment definitions). Compared to using image-only observations, this reduces training time and computational cost while allowing us to demonstrate the effectiveness of our reward function. We consider the subset of tasks that can be reliably solved using the dense Meta-World rewards, which are manually specified for each environment based on the full state of the environment. Specifically, we include tasks that can be solved with $\geq$98% success within a maximum training length of 800,000 steps. We further split these in half into 12 easy (learned in <<240,000 steps) and 13 hard tasks (240k – 800k steps).
@@
sparse only VLC MW50 VLC Open X
easy (12) 53 $\pm$ 10 65 $\pm$ 9 56 $\pm$ 11
hard (13) 62 $\pm$ 8 91 $\pm$ 4 73 $\pm$ 9
mean 58 $\pm$ 6 78 $\pm$ 5 65 $\pm$ 7
@@ 4.1 Multi-task reward function
The policy training results are shown in Table 2, with learning curves for the hard task set in Fig. 2. To summarize learning speed with a single number, we report success rate of the policy evaluated at the training length at which the manually specified Meta-World reward solves the task to $\geq$98% success. VLC trained on MW50 enables improved sample efficiency relative to the sparse reward only, which demonstrates that VLC can sufficiently generalize to trajectories not seen in demonstration data, and can effectively represent task progress for multiple tasks at once. However, a few tasks, such as Handle Press, are learned in so few trials even with sparse reward alone that there is little room for improvement in reward design, and learning is instead bottlenecked by the policy training’s sample efficiency. This is why the biggest gains are obtained for the harder tasks.
@@ 4.2 Task generalization to unseen environments
Success rates of RL policies on these 6 held-out tasks are shown in Table 3 (calculated similarly to Table 2), with learning curves in Fig. 3. In most tasks, the addition of VLC rewards improves sample efficiency of RL training despite the tasks having never been seen in reward function training, with average success rates 31 percentage points over the sparse baseline and an average sample efficiency improvement of over 3x.
In addition to task generalization, we further evaluate VLC in out-of-distribution visual conditions in Appendix F.
@@
LOReL RoboCLIP LIV Voltron R3M VLC (ours) sparse only
easy (2) 56 $\pm$ 37 46 $\pm$ 45 30 $\pm$ 18 18 $\pm$ 18 22 $\pm$ 12 57 $\pm$ 37 39 $\pm$ 25
hard (4) 59 $\pm$ 18 51 $\pm$ 14 44 $\pm$ 19 35 $\pm$ 14 24 $\pm$ 18 80 $\pm$ 10 42 $\pm$ 18
mean 58 $\pm$ 11 49 $\pm$ 15 39 $\pm$ 13 29 $\pm$ 11 16 $\pm$ 12 72 $\pm$ 13 41 $\pm$ 18
@@ 4.3 Embodiment generalization to unseen domains
The advantage of our method, and pretraining a reward function in general, is that no data collection on the target robot and in the target environment is required. To demonstrate this, we train VLC on cross-embodiment data from Open X-Embodiment (Open X-Embodiment Collaboration et al., 2023). We use the language-annotated subset, with a total of 698,000 episodes of diverse tasks filmed in various real-world robotic labs. Although some of this data does feature the Sawyer robot used in Meta-World simulations, this is only a marginal subset of 0.33% of the language-annotated videos. Moreover, the domain gap remains significant due to real-world variations in objects, backgrounds, lighting conditions, task instances and instruction formats, as well as the embodiment gap between the simulated and the real robots.
We successfully train policies (see Table 2 and Fig. 2) using Open X trained models despite a significant domain gap, highlighting the generalizability of large-scale vision-language training. We obtain an average 2x sample efficiency gain over the sparse reward in tasks solved by both rewards (sample efficiency is ill-defined if either does not solve the task), with particularly large improvements in Handle Pull Side (7x), Reach Wall (7x) and Slide Plate (6x), and an average 7 percentage point success rate increase across all 25 tasks despite misrepresenting a few tasks. Note that unlike RT-X (Open X-Embodiment Collaboration et al., 2023), our method does not use action labels, and remains equally applicable on observation-only data.
@@ 4.4 Comparison to prior work
To validate VLC’s benefits, we compare its performance to prior language-conditioned reward models LOReL (Nair et al., 2022a), RoboCLIP (Sontakke et al., 2023), LIV (Ma et al., 2023a), Voltron (Karamcheti et al., 2023) and R3M (Nair et al., 2022b), each fine-tuned on MW40. A breakdown of the key differences between VLC and these baseline methods is shown in Table 6 in Appendix E; in summary: VLC is the only method to use a sequence ranking objective or a temporal aggregation Transformer, and one of only two to use history conditioning. For the LOReL baseline, we used the proposed binary classification objective, reversed negatives and 2-frame conditioning (first and last), while keeping the architecture and CLIP pretraining identical to our method. This is to ensure LOReL’s smaller and older original architecture and lack of visual representation pretraining did not account for any difference in performance. The original implementations were used for all other methods. Further training and implementation details for the baselines are also included in the Appendix E.
We find VLC’s combination of cross-entropy and the sequential ranking objective, temporal Transformer architecture as well as full video conditioning to produce more informative reward predictions than existing methods, as shown by faster policy training on average in Table 3 and Fig. 3. For increased statistical significance, we use 5 additional random seeds (10 total) for VLC and the two strongest baselines: RoboCLIP and LOReL. Moreover, on qualitative inspection of the shape of the predicted rewards (Appendix G), we find VLC’s outputs to better distinguish successes from failures compared to either RoboCLIP or LOReL. Thanks to its conditioning on more frames of execution history and the sequential loss term, VLC also produces rewards that more smoothly increase over successful episodes than either prior method.
@@ 4.5 Model-based evaluation
We find VLC to generalize well to these tasks zero-shot, with 80% mean accuracy on held-out tasks, outperforming RoboCLIP (46%) and LOReL (36%). A breakdown per task is shown in Table 3.
5 Limitations and future work
Suboptimal demonstrations
Our sequence ranking objective most benefits from optimal or high-quality trajectories. However, this is also true for imitation learning approaches, and learning a reward function (unlike imitation learning) still allows the policy to outperform the existing trajectories by executing the task faster or by finding shortcuts: the highest rewards are still assigned to the final states, so getting to one faster is encouraged. Moreover, our objective does not impose consistent increase, only non-decrease of rewards. Therefore, a suboptimal section of a demonstration can be assigned a constant reward, representing non-progress in the task, and the objective only assumes that the suboptimal parts do not make negative progress in the task, which is a much less restrictive assumption.
Choice of evaluation benchmarks
In our formulation, we focused on tasks whose rewards can be simplified to goal reaching, but other rewards may be relevant (such as penalizing large actions). We further assumed that a sparse task completion reward can be provided, which may not always be the case.
In this work, we focused on simulated robotic manipulation tasks, and evaluation on a real robot is left for further work. Moreover, extending VLC to learn from videos of humans performing manipulation tasks, such as from the Something-Something dataset (Goyal et al., 2017), is a promising avenue for pushing the generalization capabilities of pretrained reward functions in the future.
@@ 6 Conclusion
We proposed Video-Language Critic (VLC), a method for training a foundation reward model for vision-language manipulation. In particular, we train our critic using contrastive video-language alignment and a ranking loss encouraging monotonic increases for successful trajectories. Our model predicts a language-conditioned state-value function conditioned on only a history of image observations, and can therefore readily be scaled to leverage external observation data from other actors. VLC can be used for various downstream tasks, such as model-free and model-based reinforcement learning. Further, we experimentally validated its usefulness as a reward function in out-of-distribution tasks (unseen, held-out environments in the same domain) and out-of-domain tasks (an unseen environment and embodiment in a different domain). Our experiments on Meta-World demonstrated improved results compared to 5 prior methods and a sparse reward baseline, with success rate increases of 14 and 31 percentage units, respectively, and sample efficiency gains of 30% and 300%, respectively. Unlike methods based on predicting expert actions in a given embodiment, VLC remains applicable in the absence of action labels. Therefore, it can be extended to equally learn from videos of humans in future work.
@@
Appendix A Model ablations on VLMbench
@@ A.1 VLMbench dataset
We use the VLMbench manipulation task suite to develop and validate our method without any Meta-World specific tuning. For this purpose, we collect 2700 video demonstrations and 1600 failure cases from variations of the picking task – covering different object shapes, sizes, colors and relative positions, as well as distractor objects. The natural language instructions match this diversity in task variants, such as Grasp the cylinder or Grasp the cyan object, and require distinguishing relevant objects from distractors with either absolute (color, shape) or relative (size, position) properties, such as the the larger or the front object. For more details on the benchmark, see Zheng et al. (2022).
We use these VLMbench videos to validate VLC design decisions, but defining a single informative metric on the dataset of video-caption pairs $(v^{i},c^{i})$, i=1,,Ni=1,...,N, is difficult. Test loss, video retrieval metrics such as mean recall, or classification metrics such as area under the ROC curve do not correspond well to the models’ ability to model task progress. The main difficulty is that part of the caption-to-video matching task can be solved by simply connecting objects referred to in the caption to objects present in the scene, without considering temporal information or actual task success.
@@ A.2 Ablation results
We compare two temporal aggregation methods as proposed by Luo et al. (2022): the sequence transformer and the tight-type transformer. The sequence transformer aggregates the sequence of image features $\text{ViT}(v_{1}),\text{ViT}(v_{2}),...,\text{ViT}(v_{T})]$ into a single embedding vector $S_{\theta}(v_{1:T})$, which it then compares to the caption embedding $\text{TextEnc}(c)$ with cosine similarity. The tight-type transformer, on the other hand, includes the caption embedding as an additional input to the temporal aggregator $S_{\theta}(v_{1:T},c)$, as shown in Fig. 1.
We report the results of our ablation study in Table 4. In addition to the choice of architecture, we observe performance gains from adding image augmentations from the Albumentations library (Buslaev et al., 2020), by sampling frames randomly from uniform intervals instead of deterministic uniform sampling, the addition of the sequence ranking term, as well as considering failure examples only as negatives in the contrastive objective, and report results using these settings in Section 4.
@@
Architecture $\alpha$ Data augmentations Failures Accuracy (%)
@@
Appendix B Model ablations on Meta-World
In order to isolate the contributions of the VLC architecture and of the sequence ranking objective to the observed performance improvements on Meta-World, we run an additional ablation using only the contrastive component of the loss (by setting $\alpha$ = 0). Moreover, as a second ablation variant, we use our architecture together with the LIV objective.
The resulting training curves are shown in Figure 4, with the success rates included in Table 4. All policy training runs were repeated with 10 random seeds. The full VLC objective outperforms both ablations on average across the 6 Meta-World held-out tasks – by 5 and by 12 percentage points, respectively. Although the no-sequence-ranking baseline obtained the highest success rate on the easy tasks, as there are only two tasks in this set, results on the intermediate set and the average across all tasks are more meaningful. The gains over the no-sequence-ranking ablation are small but consistent with the gains observed on VLMbench data, our second evaluation benchmark (see Appendix A). Furthermore, it is possible that tuning the $\alpha$ parameter on Meta-World would improve performance, but we did not explore this and instead tuned all hyperparameters on VLMbench only.
Note that applying the LIV loss to our architecture requires the time-contrastive loss on images to be applied after the visual encoder but before the temporal aggregator, whereas the vision-language contrastive loss is applied to the final output.
Figure 4: Ablations on the choice of objective on held-out Meta-World tasks, compared to the full VLC objective. $\alpha=0$ refers to only using the contrastive objective, while the LIV loss is a combination of the contrastive objective and the VIP (value-implicit pretraining; (Ma et al., 2023b)) loss applied on the image representations. Both ablations use the VLC architecture conditioned on 12 frames.
Table 5: Success rates (%) of VLC and ablations trained on MW40, evaluated on 6 unseen tasks (2 easy, 4 hard) at the training length at which the Meta-World hand-designed reward solves the task.
VLC VLC ($\alpha$ = 0) LIV loss, VLC archit. sparse only
easy (2) 57 $\pm$ 37 70 $\pm$ 23 46 $\pm$ 35 39 $\pm$ 25
hard (4) 80 $\pm$ 10 66 $\pm$ 19 67 $\pm$ 19 42 $\pm$ 18
mean 72 $\pm$ 13 67 $\pm$ 18 60 $\pm$ 12 41 $\pm$ 18
Appendix C Training details
C.1 Reward training